OpenStudy (anonymous):

will give medal & fan! 3 more to go..

OpenStudy (anonymous):

@Chelsea04

OpenStudy (anonymous):

do you know the circle equation?

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2=r^2\] (h,k) is the centre and r is the radius

OpenStudy (anonymous):

okay! so what do I do with 14?

OpenStudy (anonymous):

that is the radius, so if the radius is 14, that means r=14

OpenStudy (anonymous):

also, you need your centres to put into the equation

OpenStudy (anonymous):

Ohhh so what are the centres? (:

OpenStudy (anonymous):

*centre (no plural.it's a circle) The question said that the centre (of the circle) is at the origin

OpenStudy (anonymous):

so (h,k) refers to this centre. h is the x-coordinate for the centre k is the y-coordinate for the centre

OpenStudy (anonymous):

oh alrighty! so how do you find it?

OpenStudy (anonymous):

origin (from the question)

OpenStudy (anonymous):

the origin is 14?(:

OpenStudy (anonymous):

no the origin is (0,0) coordinate point

OpenStudy (anonymous):

think of the cartesian plane: at the centre where the two lines meet it is called the origin. we also know that here the x-co is 0 and the y-co is 0 therefore the origin is (0,0)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so...if h=0 and k=0 and r=14, put those values into the equation I gave you earlier and voila!

OpenStudy (anonymous):

Oh! 14?(:

OpenStudy (anonymous):

@Chelsea04

OpenStudy (anonymous):

no, the equation is: \[(x-h)^2+(y-k)^2=r^2\] so, \[(x-0)^2+(y-0)^2=14^2\] now simplify :)

OpenStudy (anonymous):

but it's 0 :(

OpenStudy (anonymous):

so... 0^2 is just 0 that means the eqn is: x^2+y^2=196 done!

OpenStudy (anonymous):

you're looking for an equation, not a value :)

OpenStudy (anonymous):

oHHHHHHH Lol! Thank youuu I have another radius question! 2nd to last one yay